# Does Time Have Its Place? Temporal Heads: Where Language Models Recall Time-specific Information

### v1

@@ Abstract
While the ability of language models to elicit facts has been widely investigated, how they handle temporally changing facts remains underexplored. We discover Temporal Heads, specific attention heads primarily responsible for processing temporal knowledge through circuit analysis. We confirm that these heads are present across multiple models, though their specific locations may vary, and their responses differ depending on the type of knowledge and its corresponding years. Disabling these heads degrades the model’s ability to recall time-specific knowledge while maintaining its general capabilities without compromising time-invariant and question-answering performances. Moreover, the heads are activated not only numeric conditions (“In 2004”) but also textual aliases (“In the year …"), indicating that they encode a temporal dimension beyond simple numerical representation. Furthermore, we expand the potential of our findings by demonstrating how temporal knowledge can be edited by adjusting the values of these heads [^1] Our datasets and code are publicly available at https://github.com/dmis-lab/TemporalHead.
@@
Yein Park1, Chanwoong Yoon1, Jungwoo Park1,3, Minbyul Jeong2thanks: Corresponding authors, Jaewoo Kang1,311footnotemark: 1 Korea University1  Upstage AI2  AIGEN Sciences3 {522yein, cwyoon99, jungwoo-park, kangj}@korea.ac.kr
@@ 1 Introduction
This profound and intricate relationship between memory and truth resonates deeply with one of the central challenges in modern artificial intelligence. While large language models (LLMs) like GPTs (OpenAI, 2022, 2024a, 2024b) and LLaMA familiess (Touvron et al., 2023a, b; Dubey et al., 2024) have demonstrated remarkable capabilities in leveraging factual knowledge, they face a unique challenge that mirrors human memory: the accurate representation of temporal knowledge—facts that transform across different time points.
@@
Figure 2: Overview of temporal knowledge circuit analysis. (A): Construct temporal knowledge circuits (TKCs), and compare it with general knowledge circuits (KCs) using time-invariant knowledge. Circuits reproduce residual streams for time T, subject S and relation R. This verifies temporal heads only found in each different TKCs of various year $T_{k}$. (B): Example of simplified TKC. Here, basic knowledge nodes is colored violet, (common in both), while Temporal Heads is highlited. (C): Attention map for temporal heads. a15.h0 means the 15th layer’s first attention head. Each head’s attention pattern is represented as the output logits of the hean by mapping to vocabulary space. Queries are input tokens focusing on others, while keys are the tokens being focused on. Values represent attention weights, indicating the strength of this focus. Total results are in Figures 78 and 911.
@@
To address them, we apply Circuit Analysis (Elhage et al., 2021; Wang et al., 2023) to reconstruct the model’s computations via localized subgraphs of attention heads, feed-forward networks (FFN), and residual streams. Especially, by systematic ablating (zeroing out) attention heads or MLP components, it pinpoints which parts are responsible for eliciting knowledge in each knowledge recalling inference (Yao et al., 2024). These knowledge circuits enable to measure how much each nodes or edges in subgraph contribute to processing facts.
@@
One step further, we explore in-depth impacts of temporal heads among different years, knowledge and conditioning types. Ablating them exclusively affects temporal information, while having negligible impact on time-invariant knowledge and general QA performance. Notably, these temporal heads are activated for both numerical expressions (“In 2004”) and textual conditions (“In the year the Summer Olympics were held in Athens”), indicating that they encode a broader temporal dimension beyond simple numerical representation. Building on this, we demonstrate that temporal knowledge editing-selectively adding their activations-enables direct intervention in year-conditioned factual recall. Through this targeted manipulation, our experiments demonstrate that the temporal heads serve as key subcomponents for encoding and modifying time-sensitive knowledge.
@@ 2.2 Knowledge Circuit
Unlike a generic circuit for any functionality, a knowledge circuit specifically captures the local subgraph dedicated to storing and relaying factual content for the knowledge triplet at hand. We specifically utilize effective attribution pruning-integrated gradients (EAP-IG), which ablating (zeroing) candidate edges and measuring drops in correct prediction (Hanna et al., 2024). For more details, see Appendix 8.1.
@@ 3 Knowledge Circuit Deciphers Temporal Head in LLMs
We now explore how knowledge circuits, extracted via EAP-IG pruning, can reveal specialized Temporal Heads in large language models (LLMs). We extend knowledge circuits in 2.2 to temporal knowledge circuits by analyzing how the same subject–relation pair can produce different objects across multiple time points. Specifically, we seek to identify which edges encode time-dependent specificity, such that an edge $e_{i}$ is crucial for predicting the time-relevant object $o_{k}$ at period $T_{k}$. Given a knowledge circuit score $S(e_{i})$ (Eq. 3), we define its temporal variant as follows:
@@
where $T_{k}$ indicates a specific time (or period), and $o_{k}$ is the corresponding object for subject $s$ and relation $r$ at time $T_{k}$. Thus, $S(e_{i},T_{k})$ measures the contribution of edge $e_{i}$ to correctly predicting $o_{k}$ under time $T_{k}$. For highlighting importance and simplifying graphs, edges retained in the temporal circuit satisfy $S(e_{i},T_{k})>\tau$, ensuring they encode time-dependent knowledge. Here, we decide to attach temporal conditioning in front of subject, following prior insight from causal tracing (§8.2) and details in Appendix 8.3.
@@ 3.1 Implementations
We conduct experiments primarily on three LLMs: Llama-2-7b-chat-hf (Touvron et al., 2023b), Qwen1.5-7B-Chat (Bai et al., 2023; Team, 2024), Phi-3-mini-4k-instruct (Abdin et al., 2024). We adopt transformer lens (Nanda and Bloom, 2022) to intercept and ablate model components, enabling EAP-IG-based circuit discovery. We mainly illustrate results on Llama2, though similar trends emerge in the other models. More details are described in Appendix 8.1.1.
@@ 3.1.1 Circuit Reproduction Score
where $d=\max\{B,0\}$. The factor $\sigma\in(0,1]$ accounts for sign mismatches, adjusting for cases where the pruned circuit’s output deviates in direction from the full model. A higher CRS indicates better reproduction of the full model’s predictions. Details on hyperparameters and adjustments are deferred to the Appendix 8.4.
@@ 3.2 Dataset
Our dataset comprises (statistics in Appendix 8.5):
@@ 3.3 Evaluation
After pruning less-contributory nodes via EAP-IG, we measure how well the resulting subgraph preserves the model’s original performance on each knowledge type. Table 1 and 34 shows the average number of nodes and edges in these pruned circuits, along with their CRS. We then apply threshold $\tau$ to remove edges/nodes that contribute marginally to object prediction, retaining only edges with scores above $\tau$ and their corresponding nodes.
@@ 3.4 Findings
What is impressive stood out in the attention heads. Temporal Heads, appearing in almost every temporal knowledge circuits but not time invariants, are shown: a15.h0, a18.h3 in Llama2. Those temporal heads reoccur across multiple year-specific circuits, and it is different for other model’s cases like a17.h15 for Qwen 1.5 in Table 2. Visualizing their attention maps in Figure 2 (C) indicates a strong focus on “In 19xx” and subsequent subject phrases, as key tokens revolve around temporal conditions with queries hooking into the subjects. This pattern corroborates the idea that these heads facilitate year-subject binding—justifying the label “temporal”, as this kind of task specific attention heads were previously suggested by Wang et al., 2023; Merullo et al., 2024; Chughtai et al., 2024; Wu et al., 2024 and  Zheng et al., 2024.
@@ 4 In-Depth Analysis of Temporal Heads
Ablation 25.6 $\Downarrow$ 61.7 $\downarrow$ 54.9 $\downarrow$
@@
Ablation 19.8 $\Downarrow$ 62.6 $\downarrow$ 49.5 $\downarrow$
@@ 4.1 Attention Head Ablation Inference
where $p_{\text{ablate}}$ denotes the log-probability computed by forward pass of model, ablating corresponding heads. This evaluates how the probability distribution over $O$ shifts, rather than just predicting the most likely answer. Details in Appendix 8.6.
@@ 4.1.2 Result of Time Invariant Knowledge
By contrast, ablating the same heads for invariant knowledge (e.g., fruit inside color) causes minimal performance drop in Table 2 and Figure 4. This indicates that “temporal heads” indeed route only temporally conditioned knowledge, and disabling them forces the model to make temporally incorrect rather than incorrect of stable knowledge. Besides, Phi-3-mini affects more sensitively than others as its parameter size is half of other two models, resulting more reactive to small changes in attention alignment. This even causes a slight gain of performance in time-invariant knowledge tasks.
@@ 4.1.3 Result of General QA
As Table 2 and result in Appendix 8.7 shows, ablating temporal heads doesn’t harm common knowledge recalling or answering general knowledge questions. Here, we test TriviaQA and Math ChroKnowledge and find out that just ablating temporal heads doesn’t effect the performance of basic QA, droping almost less than 0.6 in f1 score.
@@ 5 Temporal Knowledge Editing
Details of adding an attention is in Appendix 8.8.
@@ 6.1 Temporal Knowledge of LLM
Despite advancements in LLMs, handling temporal knowledge remains a key challenge. While prior works focus on factual consistency (Petroni et al., 2019; Kassner and Schütze, 2020) or refining model editing in MLP layer (Mitchell et al., ; Meng et al., 2022, ), few address how facts evolve over time. Studies on time-aware QA and temporal probing (Chen et al., 2021; Zhang and Choi, 2021; Dhingra et al., 2022; Jang et al., 2022) reveal that LLMs struggle with dynamically shifting facts. Recent approaches attempt explicit temporal alignment (Kim et al., 2024; Zhao et al., 2024; Mousavi et al., 2025; Park et al., 2025), but have focused on external evaluations. Our findings highlight that LLMs encode temporal facts implicitly, relying on manipulable attention heads, underscoring the need for better temporal supervision and disentangled knowledge representations.
@@ 7 Conclusion
While our approach demonstrates promising results in identifying and analyzing temporal knowledge circuits, we acknowledge some limitations in our current work. First, analysis of unstructured temporal QAs like General ChroKnowledge (Park et al., 2025) were constrained, as the underlying multiple-choice options in those tasks typically do not exhibit temporal dependencies. So we focused more on our temporal knowledge dataset, abundantly describing the effect of ablation in these cases.

### v2

@@ Abstract
While the ability of language models to elicit facts has been widely investigated, how they handle temporally changing facts remains underexplored. We discover Temporal Heads, specific attention heads that primarily handle temporal knowledge, through circuit analysis. We confirm that these heads are present across multiple models, though their specific locations may vary, and their responses differ depending on the type of knowledge and its corresponding years. Disabling these heads degrades the model’s ability to recall time-specific knowledge while maintaining its general capabilities without compromising time-invariant and question-answering performances. Moreover, the heads are activated not only numeric conditions (“In 2004”) but also textual aliases (“In the year …”), indicating that they encode a temporal dimension beyond simple numerical representation. Furthermore, we expand the potential of our findings by demonstrating how temporal knowledge can be edited by adjusting the values of these heads [^1] Our datasets and code are publicly available at https://github.com/dmis-lab/TemporalHead.
@@
Yein Park1, Chanwoong Yoon1, Jungwoo Park1,3, Minbyul Jeong2thanks: Corresponding authors, Jaewoo Kang1,311footnotemark: 1 1Korea University  2Upstage AI  3AIGEN Sciences {522yein, cwyoon99, jungwoo-park, kangj}@korea.ac.kr
@@ 1 Introduction
This profound and intricate relationship between memory and truth resonates deeply with one of the central challenges in modern artificial intelligence. While large language models (LLMs) like GPTs (OpenAI, 2022, 2024a, 2024b) and LLaMA families (Touvron et al., 2023a, b; Dubey et al., 2024) have demonstrated remarkable capabilities in leveraging factual knowledge, they face a unique challenge that mirrors human memory: the accurate representation of temporal knowledge—facts that transform across different time points.
@@
Figure 2: Overview of temporal knowledge circuit analysis. (A): Construct temporal knowledge circuits (TKCs), and compare it with general knowledge circuits (KCs) using time-invariant knowledge. Circuits reproduce residual streams for time T, subject S and relation R. This verifies temporal heads only found in each different TKCs of various year $T_{k}$. (B): Example of simplified TKC. Here, basic knowledge nodes is colored violet, (common in both), while Temporal Heads is highlighted. (C): Attention map for Temporal Heads. a15.h0 means the 15th layer’s first attention head. Each head’s attention pattern is visualized with the attention weight assigned by the queries (row) to the keys (column). Queries are the tokens distributing attention, and Keys are the tokens receiving attention. Values represent attention weights, indicating the strength of this focus. Total results are in Figures 711.
@@
To address them, we apply Circuit Analysis (Elhage et al., 2021; Wang et al., 2023) to reconstruct the model’s computations via localized subgraphs of attention heads, feed-forward networks, and residual streams. Especially, by systematic ablating (zeroing out) attention heads or multilayer perceptron (MLP) components, it pinpoints which parts are responsible for eliciting knowledge in each recalling tasks (Yao et al., 2024). These knowledge circuits enable to measure how much each nodes or edges in subgraph contribute to processing facts.
@@
One step further, we explore in-depth impacts of temporal heads among different years, knowledge and conditioning types. Ablating them exclusively affects temporal information, while having negligible impact on time-invariant knowledge and general question answering performance. Notably, these temporal heads are activated for both numerical expressions (“In 2004”) and textual conditions (“In the year the Summer Olympics were held in Athens”), indicating that they encode a broader temporal dimension beyond simple numerical representation. Building on this, we present that temporal knowledge editing-selectively adding their activations-enables direct intervention in year-conditioned factual recall. Through this targeted manipulation, our experiments demonstrate that the temporal heads serve as key subcomponents for encoding and modifying time-sensitive knowledge.
@@ 2.2 Knowledge Circuit
Unlike a generic circuit for any functionality, a knowledge circuit specifically captures the local subgraph dedicated to storing and relaying factual content for the knowledge triplet at hand. We specifically utilize effective attribution pruning-integrated gradients (EAP-IG), which ablating (zeroing) candidate edges and measuring drops in correct prediction (Hanna et al., 2024). For more details, see Appendix A.
@@ 3 Knowledge Circuit Deciphers Temporal Head in LLMs
We now explore how knowledge circuits, extracted via EAP-IG pruning, can reveal specialized Temporal Heads in large language models (LLMs). We extend knowledge circuits in §2.2 to temporal knowledge circuits by analyzing how the same subject–relation pair can produce different objects across multiple time points. Specifically, we seek to identify which edges encode time-dependent specificity, such that an edge $e_{i}$ is crucial for predicting the time-relevant object $o_{k}$ at period $T_{k}$. Given a knowledge circuit score $S(e_{i})$ (Eq. 3), we define its temporal variant as follows:
@@
where $T_{k}$ indicates a specific time (or period), and $o_{k}$ is the corresponding object for subject $s$ and relation $r$ at time $T_{k}$. Thus, $S(e_{i},T_{k})$ measures the contribution of edge $e_{i}$ to correctly predicting $o_{k}$ under time $T_{k}$. For highlighting importance and simplifying graphs, edges retained in the temporal circuit satisfy $S(e_{i},T_{k})>\tau$, ensuring they encode time-dependent knowledge. Here, we decide to attach temporal conditioning in front of subject, following prior insight from causal tracing (§B) and details in Appendix C.
@@ 3.1 Implementations
We conduct experiments primarily on three LLMs: Llama-2-7b-chat-hf (Touvron et al., 2023b), Qwen1.5-7B-Chat (Bai et al., 2023; Team, 2024), Phi-3-mini-4k-instruct (Abdin et al., 2024). We adopt transformer lens (Nanda and Bloom, 2022) to intercept and ablate model components, enabling EAP-IG-based circuit discovery. We mainly illustrate results on Llama2, though similar trends emerge in the other models. More details are described in Appendix A.1.
@@ 3.1.1 Circuit Reproduction Score
where $d=\max\{B,0\}$. The factor $\sigma\in(0,1]$ accounts for sign mismatches, adjusting for cases where the pruned circuit’s output deviates in direction from the full model. A higher CRS indicates better reproduction of the full model’s predictions. We describe the details of hyperparameters and adjustments to the Appendix D.
@@ 3.2 Dataset
Our dataset comprises (statistics in Appendix E):
@@ 3.3 Evaluation
After pruning less-contributory nodes via EAP-IG, we measure how well the resulting subgraph preserves the model’s original performance on each knowledge type. Table 1 and 45 show the average number of nodes and edges in these pruned circuits, along with their CRS. We then apply threshold $\tau$ to remove edges/nodes that contribute marginally to object prediction, retaining only edges with scores above $\tau$ and their corresponding nodes.
@@ 3.4 Findings
What stood out most was found in the attention heads. Temporal Heads, appearing in almost every temporal knowledge circuits but not time invariants, are shown: a15.h0, a18.h3 in Llama2. Those temporal heads reoccur across multiple year-specific circuits, and it is different for other model’s cases like a17.h15 for Qwen 1.5 in Table 2. Visualizing their attention maps in Figure 2 (C) indicates a strong focus on “In 19xx” and subsequent subject phrases, as key tokens revolve around temporal conditions with queries hooking into the subjects. This pattern corroborates the idea that these heads facilitate year-subject binding—justifying the label “temporal”, as this kind of task specific attention heads were previously suggested by Wang et al., 2023; Merullo et al., 2024; Chughtai et al., 2024; Wu et al., 2024 and  Zheng et al., 2024.
@@ 4 In-Depth Analysis of Temporal Heads
Ablation 25.6 $\Downarrow$ 61.7 54.9
@@
Ablation 19.8 $\Downarrow$ 62.6 49.5
@@ 4.1 Attention Head Ablation Inference
where $p_{\text{ablate}}$ denotes the log-probability computed by forward pass of model, ablating corresponding heads. This evaluates how the probability distribution over $O$ shifts, rather than just predicting the most likely answer. Details in Appendix F.
@@ 4.1.2 Result of Time Invariant Knowledge
By contrast, ablating the same heads for invariant knowledge (e.g., fruit inside color) causes minimal performance drop in Table 2 and Figure 4. This indicates that “temporal heads” indeed route only temporally conditioned knowledge, and disabling them forces the model to make temporally incorrect rather than incorrect of stable knowledge. Besides, Phi-3-mini-4k-instruct affects more sensitively than others as its parameter size is half of other two models, resulting more reactive to small changes in attention alignment. This even causes a slight gain of performance in time-invariant knowledge tasks.
@@ 4.1.3 Result of General QA
As Table 2 and result in Appendix G shown that ablating temporal heads doesn’t harm common knowledge recalling or answering general knowledge questions. Here, we test TriviaQA and Math ChroKnowledge and find out that just ablating temporal heads doesn’t affect the performance of basic QA, droping almost less than 0.6 in F1 score.
@@ 4.2 Alias Test With Textual Conditioning
Set Temporal Knowledge (%) Avg
Spo Prez CEO Def Mov GDP Infla
Fundamental Prompt: In XXXX, Lionel Messi was a member of …
Base 41.9 80.7 27.5 13.5 23.1 10.4 10.8 29.7
Abl 40.0 75.6 21.3 13.3 9.37 10.7 9.34 25.6
Prompt Variation: In year XXXX, Lionel Messi was a member of …
Base 40.5 82.6 45.8 13.6 22.1 17.3 12.1 33.4
Abl 39.7 75.4 43.2 13.2 14.5 14.1 10.3 30.0
Real-World Question: In XXXX. which sports team was Lionel …
Base 40.7 81.5 55.6 10.1 24.1 19.2 10.5 34.5
Abl 40.5 74.8 42.6 10.1 16.4 17.8 8.76 30.1
Table 3: Results of prompt variations, comparing baseline performance to the ablated model (i.e., removing Temporal Heads). Categories of temporal knowledge are same as Table 8, which is Sports, Presidents, CEO, Defense, Movies, GDP and Inflations. Each scores were measured in probability (%) with averaging effect of multiple heads ablation results (a15.h0 and a18.h3 for Llama2). The most dropped score for each column is colored red.
4.3 Additional Test With Prompt Variations
For more qualitative experiments, particularly regarding prompt settings, we do more analysis study with various prompt styles. As we originally used only one fundamental prompt format (e.g., "In XXXX"), we have conducted additional ablation experiments with Llama2 under new prompt settings to address concerns about generalizabilty.
  • Variation of Fundamental Prompt maintains the core temporal format while adding slight textual variation (“In year XXXX”).
  • Real-World Question Format simulates a more practical Q&A scenario, moving beyond a declarative statement into a direct question.
We evaluate each prompt settings on multiple knowledge categories in Table 3 same as Table 8 in Appendix. Here, we find out key observations. Across all cases, ablating Temporal Heads (i.e., disabling them) yields a notable drop in performance for each knowledge category. So, even when the prompt style changes from the original “In XXXX” to “In year XXXX” or a question-based format, the importance of Temporal Heads remains evident. These results reinforce our claim that Temporal Heads underlie the model’s ability to handle time-sensitive knowledge, regardless of prompt diversity, demonstrating our approach’s generalizability.
4.4 Difference with General Formats
Interestingly, if we suggest prompt without explicit temporal aspects such as "Who is the president of South Korea?" and "The president of Argentina is", the results indicated no meaningful nodes or edges in every circuits. Using Llama2, we conduct two approaches to addressing these scenarios:
  • Construct circuits by comparing target object with objects from different years as corrupted run (Isolating only the temporal component).
  • Construct circuits by comparing target object with objects from different knowledge contexts as corrupted run (e.g., presidents of different countries).
In each results, CRS dropped significantly (e.g., 61 to 18 in same knowledge category), reflecting weak or absent circuit reconstruction in scenarios lacking explicit temporal conditions. Still, ablating Temporal Heads affected the model’s greedy decoding outputs (e.g., switching the answer from "Moon Jae-in" to a temporally alternative object in different year "Park Geun-hye" in response to prompt “The president of South Korea is”). This implies that such queries, despite appearing static, inherently involve subtle temporal components. Thus, we conclude that it is challenging to construct temporal knowledge circuit without temporal conditions as it is inherently temporal, and it is hard to distinguish model’s activation difference between objects without temporal aspects, which needs to construct circuits. This finding reinforces the validity of our methodology—isolating the unique characteristics of temporal knowledge through comparison with commonsense knowledge.
@@ 5 Temporal Knowledge Editing
Details of adding an attention is in Appendix H.
@@
Furthermore, we applied attention addition-based temporal knowledge editing to those prompt variations in Section 4.3. Although the number of successful edited cases decreased approximately by half in the Llama2 model compared to the fundamental setting (from 8 to 4), the head with the highest number of success remained the Temporal Head a18.h3. This demonstrates that attention addition is also effective across diverse prompt settings, highlighting potential synergies through integration with other methodologies to further enhance temporal knowledge processing capabilities.
@@ 6.1 Temporal Knowledge of LLM
Despite advancements in LLMs, handling temporal knowledge remains a key challenge. While prior works focus on factual consistency (Petroni et al., 2019; Kassner and Schütze, 2020) or refining model editing in MLP layer (Mitchell et al., ; Meng et al., 2022, ), few address how facts evolve over time. Studies on time-aware QA and temporal probing (Chen et al., 2021; Zhang and Choi, 2021; Dhingra et al., 2022; Jang et al., 2022) reveal that LLMs struggle with dynamically shifting facts. Recent approaches attempt explicit temporal alignment (Kim et al., 2024; Zhao et al., 2024; Mousavi et al., 2025; Park et al., 2025), but have focused on external evaluations.
Beyond these approaches, interpretable studies have mathematically and empirically demonstrated that LLMs can implicitly interpolate and process continuous temporal and spatial information, as well as compositional heuristics, even without explicit context in the training corpus (implicitly; Arithmetic). Building on this, our findings highlight that LLMs encode temporal facts implicitly, relying on manipulable attention heads, underscoring the need for better temporal supervision and disentangled knowledge representations.
@@ 7 Conclusion
While our approach demonstrates promising results in identifying and analyzing temporal knowledge circuits, we acknowledge some limitations in our current work.
First, analysis of unstructured temporal QAs like General ChroKnowledge (Park et al., 2025) were constrained, as the underlying multiple-choice options in those tasks typically do not exhibit temporal dependencies. So we focused more on our temporal knowledge dataset, abundantly describing the effect of ablation in these cases. In addition, even though our approach systematically validated the importance of Temporal Heads in processing temporal knowledge, future work should include its broader application for enhancing temporal reasoning to increase practical value.